40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jun 27 23:49:56 2014, Koji, Update, General, IR beam found, TTs not aligned well yet 
    Reply  Mon Jun 30 16:55:24 2014, ericq, Update, General, IFO (mostly) aligned 
       Reply  Mon Jun 30 22:06:55 2014, Jenne, Update, ASC, IFO (mostly) aligned - ASS working 
          Reply  Wed Jul 2 16:18:32 2014, Jenne, Update, ASC, IFO aligned, PRMI + 2 arms achieved 
Message ID: 10110     Entry time: Fri Jun 27 23:49:56 2014     Reply to this: 10113
Author: Koji 
Type: Update 
Category: General 
Subject: IR beam found, TTs not aligned well yet 

The IR beam was found on the PRM surface, some CCDs, and in the X arm. The TTs are not aligned well yet.

I'm leaving the IFO with the following state.


Status:

ITMY/ETMY - aligned to the given green beam. GTRY (no PSL green) 1.0~1.1

ITMX/ETMX - aligned to the green beam. The end PZT for the green beam was steered to have maximum GTRX (0.76 without PSL green)

TT1/TT2 - unknown alignment, TT1/TT2 are related such that the spot is on the POP CCD

PRM - aligned to the given IR beam (i.e. PRM spot on the REFL CCD)

BS - aligned to the given IR beam (i.e. ITMX spot on the AS CCD, The X arm is flashing)


Notes:

- ITMX was stuck in the suspension. it was caused by the EQ.

- When the X arm was aligned to the green beam, there was no green hitting on the GTRX PD. That's why the end PZT was adjusted.

- In order to earn more range for TT1, C1:IOO-TT1_YAW_GAIN and C1:IOO-TT1_PIT_GAIN were increased to 300 (100 nominal) and the limiter (at 500) were removed.

- The HeNe laser for BS/PRM does not emit the beam even with the driver turned on. Is there a hidden shutter or something? ==> Jenne

 


TO DO:

- Find the Y arm fringe by moving TT1 and TT2 without loosing the PRM/AS/POP spots.

 

ELOG V3.1.3-